RC meeting and Run Plan, Saturday Sept. 30, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)
MCC spent all of the DAY and SWING shifts, and half of OWL, trying to address the Ion Chamber trips that was keeping Hall A from taking any CW beam. Unfortunately, this interfered with Hall C's attempts to take data, so at 4am, the decision was made to stop trying to give Hall A beam until DAY shift took over. With this, I let the SL for OWL go and get some sleep, but the TO was kept on shift to monitor alarms. Since 8 am today, the MCC has been attempting to rectify whatever problems were causing Hall A's beam issues, and this seems to have been generally successful. Beam recovery was completed, but before any production data could be taken, one of the RF sections dropped. When this is fixed, beam recovery will be attempted again, and production running will begin at 35uA.
Also, at the end of the DAY shift yesterday, Alexandre went into the hall to replace the HCAL fADC module that Jiwan attempted to look at on Thursday. This replacement seems to have been successful (https://logbooks.jlab.org/entry/4189423).
During the start of DAY shift, the newer PC controlling some of the lasers turned off. Due to the lack of laser-trained operators around, the crew had to wait for 2 and a half hours for an expert to arrive. The PC has since been restarted, and the lasers turned on.
- Whenever Hall A attempted to get CW beam on the C-hole target, the Hall A Ion Chambers would trip, and stop the beam.
- MCC spent until 4 am trying to get this fixed, calling in their Optics expert at 7:30 pm. The expert was not able to clear up the Hall A beam issues, so they left at 4 am, and MCC stopped trying to deliver beam to Hall A until this morning.
- MCC also alerted the CH to issues they saw in the Moller quads (https://logbooks.jlab.org/entry/4189495) and the beam window cooling jet air flow (https://logbooks.jlab.org/entry/4189513). These seem to either have been resolved, or to have resolved themselves, at least for now.
- MCC was able to get CW beam to Hall A at 11 am this morning (on No Target, for testing), so beam recovery was started.
- At 12:30 pm, the CH announced that they were ready to take production data on the polarized helium cell, but on of the RF segments dropped before this could be attempted. This issue has already taken over one and a half hours to address, and does not look to be close to completion.
- No maintenance day on the MCC white board (until 10/10/2023).
Communication with the PD
- The decision to stop beam to Hall A was taken at 4 am, but nothing was said to the shift crew in Hall A, nor to the RC (nor, as far as I understand, to the PD!)
- The only reason we found out about this is because the shift crew saw beam in Hall C, and called MCC to inquire about our beam.
- This communication issue is compounded by Hall A not having been told about the decision to stop beam overnight on 9/28-9/29, and only finding out second-hand after I called the MCC to ask for an update to beam recovery.
- This has been passed on to the PD, but if there are any more instances of this, let me know. This is not a polite request. I cannot complain about communication issues if I am not told about them in the first place.
Data collected :
- Accumulated corrected charge collected: 0C (no beam = no charge...) ...again, again.
- Around 30 minutes of production running was attempted after the beam was called ready, but due to DAQ issues, these have been declared as junk.
Hall A Subsystems Status
- Latest target polarization: 39.9% [12:56 pm]. This low value is due to the laser computer trip earlier in the shift, but it is not too bad for having lost lasers for 2 1/2 hours.
- Nothing has changed since yesterday.
Hall access requests
- If anyone wants to make an access, let me know! There is no planned maintenance day before Tuesday 10th, so any accesses need to be opportunistic.
- Still need to fill open slots
- Open slots on Wednesday 4th: OWL and SWING (both TO).
Please sign up for any open shifts!
Short Term Run plan
- Beam recovery to the hall
- Production run on polarized He3 target @35uA beam current.
- If the beam and target are stable at 35uA, call the RC. We can then go up to 40uA.
- If 40uA is stable for a few hours, we can consider going up to 45uA - call the RC first!
==> List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)